OpenStudy (yinshy):

What is the approximate circumference of a pizza pan with a diameter of 13.2 inches? (Use π = 3.14) A. 62.172 inches B. 20.724 inches C. 41.448 inches D. 82.896 inches

OpenStudy (kendricklamar2014):

\[C = 2 \pi r\]

OpenStudy (kendricklamar2014):

To get the radius from the diameter, all you need to do is divide the diameter by 2: \[\frac{ 13.2 }{ 2 } = ~ ?\]

OpenStudy (kendricklamar2014):

What is 13.2 / 2 @Yinshy

OpenStudy (yinshy):

um 6.6

OpenStudy (kendricklamar2014):

That is correct. Now we substitute 6.6 for r: \[C = 2 \times \pi \times 6.6\]

OpenStudy (yinshy):

so c

OpenStudy (yinshy):

lol

OpenStudy (kendricklamar2014):

Yes, that is the answer :D

OpenStudy (yinshy):

you gave me the answer lol

OpenStudy (kendricklamar2014):

How?

OpenStudy (yinshy):

the C=

OpenStudy (yinshy):

and what ever numbers

OpenStudy (kendricklamar2014):

Oh lol XD

OpenStudy (yinshy):

xD be carfull to not give the answer lol

OpenStudy (yinshy):

careful

OpenStudy (kendricklamar2014):

C is Circumference. So when I said that, I meant Circumference = 2 x pi x 6.6

OpenStudy (kendricklamar2014):

Lol :P

OpenStudy (yinshy):

ik i was just trying to be funny

OpenStudy (yinshy):

i have more questions ok

OpenStudy (kendricklamar2014):

Ik, ok ill try to help :)